​One's intelligence​ quotient, or​ IQ, varies directly as a​ person's mental age and inversely as that​ person's chronological age. A person with a mental age of 20 and a chronological age of 20 has an IQ of 120.  What is the chronological age of a person with a mental age of 35 and an IQ of 140 question mark
Answer by Boreal(15235)

IQ=k*MA/CA
120=k*20/20, so k=120
use k below
140=120*35/CA
CA=120*35/140=30 years